OpenStudy (anonymous):

By making a change of variables and letting a = 1/x and b = 1/y, the following system can be converted into a system of linear equations in terms of a and b. Solve the new system and use your solution to find the solution to the original system of equations:

OpenStudy (anonymous):

2/x - 3/y = 11/3 3/x + 2/y = -1

OpenStudy (anonymous):

^ this is the original system of equations : )

OpenStudy (anonymous):

I will walk you through with one. (6, -5) and is parallel to the line whose equation is 2x - 3y = 11? 2x - 3y = 11 add 3y to both sides 2x=11+3y -11 2x-11=3y 3y=2x-11 /3 y= (2/3)*x-11/3 slope m = 2/3 The parallel line will have the same slope as that of the given line. so slope of the required line will be 2/3 .. line passes through (6,-5) plug the value of slope , and the point (6,-5) to find b the y intercept of the required line y=mx+b -5=2/3 *6+b -5=4+b b=-9 ... so the equation has slope 2/3 and y intercept -9 y=(2/3)x-9 ...

OpenStudy (anonymous):

Oh my god thank you so much!!!!